# Generating Synthetic Electronic Health Record (EHR) Data: A Review with Benchmarking

### v1

@@ Acknowledgements
[53] EMERGE Rule 2010 - - -
[86] PeGS non-param 2013 - - -
[54] Plasmode Rule 2014 - - -
[55] PADARSER/GRiSER Rule 2014 - - -
[56] CoMSER Rule 2016 - - -
[59] RGA/RCGAN GAN 2017 - - -
[58] MedGAN GAN 2017
[89] Synthea Rule 2018 - - -
[57] ATEN Rule 2018 - - -
[60] medBGAN/medWGAN GAN 2019
[61] GcGAN GAN 2019 - - -
[22] HealthGAN GAN 2020
[62] DAAE GAN 2020
[17] EMR-WGAN/EMR-CWGAN GAN 2020 - - -
[63] HGAN GAN 2020 - - -
[64] CorGAN GAN 2020
[65] ADS-GAN GAN 2020 - - -
[66] SMOOTH-GAN GAN 2020 - - -
[87] Bayesian Network Bayesian Network 2021
[67] SynTEG GAN 2021 - - -
[18] EVA VAE 2021 - - -
[74] PromptEHR Transformer 2022
[68] LongGAN GAN 2022 - - -
[40] Graph-VAE VAE (Graph) 2023 - - -
[44] EHR-M-GAN GAN 2023 - - -
[79] EHR-TabDDPM Diffusion 2023
[80] MedDiff Diffusion 2023
[88] LSTM-EHR LSTM 2023 - - -
[81] ScoEHR Diffusion 2023
[90] HALO Transformer 2023
[82] TimeDiff Diffusion 2023
[97] EHR-Safe GAN 2023
[43] EHRDiff Diffusion 2024
[41] MSIC VAE 2024
[83] EHR-D3PM Diffusion 2024
[98] CEHR-GPT Transformer 2024 - - -
[84] FlexGen-EHR Diffusion 2024 - - -
[69] Bt-GAN GAN 2024 - - -
[42] IGAMT GAN 2024   ✓ - - -
[75] LLM-EHR Transformer/LLM 2024
[70] LDP-GAN GAN 2024 - - -
[85] EHRPD Diffusion 2024
@@
Table 2: Summary statistics of the MIMIC-III and MIMIC-IV data. For age, gender, ethnicity, and diseases, the values represent percentages; counts are shown in parentheses.
@@
(a) MIMIC-III.
(b) MIMIC-IV.
Figure 3: Boxplots of the differences in the estimated Pearson correlation for $100,000$ randomly selected pairs of phecodes (out of all (${1,773\choose 2}$ pairs) comparing the ones obtained from synthetic data with those from (a) MIMIC-III and (b) MIMIC-IV for each of the 9 selected methods. The solid red line represents an exact match in correlation; the area within the red dotted lines indicates a correlation difference of less than 0.01.
Fidelity Utility Privacy
MIMIC-III
MIMIC-IV
Obesity v.s. All Cancers Diabetes v.s. Hypertension
$\hat{\beta}$ 95% CI $\hat{\beta}$ 95% CI
Real Data MIMIC-III 0.871 (0.782 0.960) 1.219 (1.174 1.263)
MIMIC-IV 0.810 (0.782 0.838) 1.412 (1.388 1.437)
Synthetic Data Resample 0.868 (0.781 0.955) 1.194 (1.151 1.237)
Prevalence-based Random -0.125 (-0.316 0.066) 0.003 (-0.038 0.043)
Synthea 1.685 (1.614 1.756) 3.590 (3.512 3.668)
Plasmode -0.039 (-0.143 0.066) 0.027 (-0.118 0.171)
MedGAN 3.636 (3.453 3.819) 0.864 (0.822 0.905)
CorGAN 1.080 (0.970 1.191) 1.165 (1.120 1.209)
VAE - - 3.190 (3.063 3.318)
PromptEHR 0.514 (0.420 0.608) 0.597 (0.553 0.642)
EHRDiff 0.565 (0.527 0.603) 0.946 (0.891 1.000)
Figure 4: The performance under varying sizes of the synthetic data ($M$) generated by CorGAN trained on full real data. The chosen evaluation metrics for the three perspectives are Maximum Mean Discrepancy (MMD), Train on Synthetic, Test on Real (TSTR) measured by Area Under the Curve (AUC), and F1 score for Attribute Inference Risk (AIR), respectively. The synthetic data generation process is repeated five times, and the mean evaluation metrics are reported. Note that for MMD and F1 score in AIR, lower values indicate better performance.
Figure 5: The performance of the synthetic data generated by CorGAN trained under varying sizes of the real data for training ($N$). The chosen evaluation metrics for the three perspectives are Maximum Mean Discrepancy (MMD), Train on Synthetic, Test on Real (TSTR) measured by Area Under the Curve (AUC), and F1 score for Attribute Inference Risk (AIR), respectively. The synthetic data generation process is repeated five times, and the mean evaluation metrics are reported. Note that for MMD and F1 score in AIR, lower values indicate better performance.
Appendix S1 Supplementary Materials
@@
Figure S4: Boxplots of the ratio of the estimated PhecodeX-wise prevalence between the real and the synthetic data generated by CorGAN, grouped by PhecodeX categories defined in [36]. Y-axis represents the ratio of estimated prevalence, less than 1 means underestimation by the synthetic data. X-axis represents all the PhecodeX categories. Red line indicates the median of prevalence ratio of all phecodes. The significant over-estimation within the neonatal phenotype group in the MIMIC-IV dataset is because of its lack of neonatal population.
@@
Figure S7: Within-dataset Pearson correlation among different evaluation metrics in Table 3 for (a) MIMIC-III, and (b) MIMIC-IV, respectively. All “lower-is-better” evaluation metrics (e.g., Maximum Mean Discrepancy (MMD)) are flipped to their opposite numbers before calculating the Pearson correlation. Positive correlations (grids in red) indicate synergistic relationships between two metrics, and negative correlations (girds in blue) indicate trade-offs. No numbers are shown on the diagonal because here we focus on the relationships between distinct metrics evaluated on the same data; only lower triangular numbers are shown because of symmetry.

### v2

@@ 3.1 Summary of scoping review
[buczak2010data] EMERGE Rule 2010 - - -
[park2013perturbed] PeGS non-param 2013 - - -
[franklin2014plasmode] Plasmode Rule 2014 - - -
[dube2014approach] PADARSER/GRiSER Rule 2014 - - -
[mclachlan2016using] CoMSER Rule 2016 - - -
[esteban2017real] RGA/RCGAN GAN 2017 - - -
[choi2017generating] MedGAN GAN 2017
[walonoski2018synthea] Synthea Rule 2018 - - -
[mclachlan2018aten] ATEN Rule 2018 - - -
[baowaly2019synthesizing] medBGAN/medWGAN GAN 2019
[yang2019grouped] GcGAN GAN 2019 - - -
[yale2020generation] HealthGAN GAN 2020
[lee2020generating] DAAE GAN 2020
[zhang2020ensuring] EMR-WGAN/EMR-CWGAN GAN 2020 - - -
[yan2020generating] HGAN GAN 2020 - - -
[torfi2020corgan] CorGAN GAN 2020
[yoon2020anonymization] ADS-GAN GAN 2020 - - -
[rashidian2020smooth] SMOOTH-GAN GAN 2020 - - -
[kaur2021application] Bayesian Network Bayesian Network 2021
[zhang2021synteg] SynTEG GAN 2021 - - -
[biswal2021eva] EVA VAE 2021 - - -
[wang2022promptehr] PromptEHR Transformer 2022
[sun2021generating] LongGAN GAN 2022 - - -
[bing2022conditional] HealthGen VAE 2022
[SUN2023104404] DP-CGANS GAN 2023 - - -
[…]
[lu2023multi] MT-GAN GAN 2023
[nikolentzos2023synthetic] Graph-VAE VAE (Graph) 2023 - - -
[li2023generating] EHR-M-GAN GAN 2023 - - -
[ceritli2023synthesizing] EHR-TabDDPM Diffusion 2023
[he2023meddiff] MedDiff Diffusion 2023
[mosquera2023method] LSTM-EHR LSTM 2023 - - -
[naseer2023scoehr] ScoEHR Diffusion 2023
[theodorou2023synthesize] HALO Transformer 2023
[tian2023fast] TimeDiff Diffusion 2023
[yoon2023ehr] EHR-Safe GAN 2023
[yuan2023ehrdiff] EHRDiff Diffusion 2024
[sun2024collaborative] MSIC VAE 2024
[chen2024guided] EHR-D3PM Diffusion 2024
[pang2024cehr] CEHR-GPT Transformer 2024 - - -
[heflexible] FlexGen-EHR Diffusion 2024 - - -
[ramachandranpillai2024bt] Bt-GAN GAN 2024 - - -
[wang2024igamt] IGAMT GAN 2024   ✓ - - -
[vardhan2024large] LLM-EHR Transformer/LLM 2024
[gwon2024ldp] LDP-GAN GAN 2024 - - -
[zhong2024synthesizing] EHRPD Diffusion 2024
[lee2024leveraging] CodeAR VAE 2024
[karami2024timehr] TimEHR GAN 2024
  • Rule-based method: These approaches generate synthetic EHR data using predefined algorithms and rules, often incorporating real-world disease prevalence and national census data as hyperparameters to simulate disease occurrence and population demographics. From our review, 6 out of 48 methods are rule-based methods[buczak2010data, franklin2014plasmode, dube2014approach, mclachlan2016using, mclachlan2018aten].
  • GAN-based method: With the rise of deep learning, GANs have become widely adopted for synthetic EHR generation. 21 out of 48 methods are GAN-based methods[baowaly2019synthesizing, yang2019grouped, yale2020generation, lee2020generating, zhang2020ensuring, yan2020generating, torfi2020corgan, yoon2020anonymization, rashidian2020smooth, zhang2021synteg, sun2021generating, li2023generating, ramachandranpillai2024bt, wang2024igamt, gwon2024ldp].
  • VAE-based method: As another strand of deep generative modeling that are popular in the artificial intelligence (AI) community, VAEs are applied to generate synthetic EHR data from latent embeddings[biswal2021eva, bing2022conditional, lee2024leveraging]. Our review shows that 5 out of 48 methods apply VAEs to generate synthetic EHR data.
  • Transformer-based method: Transformers[vaswani2017attention], known for their ability to process sequential data[li2019neural] and their success in natural language processing[devlin2018bert, brown2020gpt3], have also been explored for longitudinal synthetic EHR generation. In our review, 5 out of 48 methods learn to generate synthetic EHR data with the transformer structures.
  • Diffusion-based method: Most recently, diffusion models[sohl2015deep, ho2020denoising], which have demonstrated strong performance in image and video generation[ho2022imagen, ho2022video, saharia2022photorealistic], have been applied for synthetic EHR generation. 8 out of the 48 studies have proposed diffusion-based models for this purpose[ceritli2023synthesizing, he2023meddiff, naseer2023scoehr, tian2023fast, yuan2023ehrdiff, chen2024guided, heflexible, zhong2024synthesizing].
In addition to the five main categories above, conventional statistical methods were also used to generate synthetic EHR data, for example, Gibbs sampling-based algorithms[park2013perturbed]; Bayesian Networks[kaur2021application]. Additionally, other neural network architectures are also explored[mosquera2023method].
Based on these literature, we identify three key functionalities that are critical for the design of synthetic data generation algorithms, which we leave in Section S2.
3.2 Methods selected for benchmarking
Despite the recent development of synthetic EHR methods, 31 out of 48 studies identified in our scoping review remain closed-source in terms of the codebases at the time of this scoping review (Table 1). To provide a benchmarking, we select seven generation methods with high impact and publicly available repositories with annotated code and detailed instruction to enable implementation on the MIMIC datasets without significantly additional efforts, chosen from the five identified methodological categories identified in our scoping review. Specifically, for rule-based methods, we select Plasmode[franklin2014plasmode] and Synthea[walonoski2018synthea]; for GAN-based methods, we choose MedGAN[choi2017generating] and CorGAN[torfi2020corgan]; for VAE-based methods, we use a baseline implemented by CorGAN’s authors due to the lack of easily reproducible codebases in existing VAE-based methods; for transformer-based methods, we select PromptEHR[wang2022promptehr]; and for diffusion-based methods, we choose EHRDiff[yuan2023ehrdiff]. For all the existing methods, we follow their default settings, except necessary minor changes for debugging.
Note that among the included methods, PromptEHR and Synthea generate EHR data longitudinally. Specifically, PromptEHR is a Transformer-based model that generates subsequent visits conditional on the sequence of prior visits; Synthea simulates full patient trajectories using disease progression modules. For evaluation, we generate data longitudinally and then aggregate it into binary phenotypes-labeling a disease as 1 if it occurs at least once. Summary statistics for the number of visits generated in the longitudinal synthetic data can be found in Table S9 in the Supplementary Materials.
Additionally, we include two simple baselines for comparison. The first baseline, Prevalence-based Random (PBR), randomly generates synthetic EHR data based on the estimated marginal phenotype prevalences without taking phenotype correlation into account. The second baseline, Resample, bootstraps the real EHR dataset to “synthesize” data. We leave the detailed description of the code preprocessing pipeline including mapping to a common coding system in Section S1.
3.3 Evaluation results
In this subsection, we introduce the benchmarking results of key evaluation metrics. Full descriptions, including the results of computational cost (Table S4) and numerical studies (Figure S10S11), sensitivity analysis on coding systems, and sub-experiment results on evaluating the methods in different sub-populations, are presented in Section S2 of the Supplementary Materials.
Fidelity
Figure 1 presents boxplots of the marginal prevalences in the real and synthetic EHR data. The PBR baseline, MedGAN, CorGAN, and EHRDiff effectively preserve the prevalence distribution in MIMIC-III, while Plasmode, Synthea, and VAE consistently underestimate it. The MIMIC-IV dataset exhibits a similar prevalence distribution to MIMIC-III, though with a slightly lower median prevalence across all phenotypes.
Figure 2 displays the number of unique phecodes per patient in EHR data. CorGAN closely matches MIMIC-III, whereas most other methods underestimate the number of unique phecodes. In contrast, PromptEHR consistently overestimates it.
Figure 3: Boxplots of the differences in the estimated Pearson correlation for $100,000$ randomly selected pairs of phecodes (out of all (${1,773\choose 2}$ pairs) comparing the ones obtained from synthetic data with those from MIMIC-III for each of the 9 selected methods. The solid red line represents an exact match in correlation; the area within the red dotted lines indicates a correlation difference of less than 0.01.
Figure 4: Boxplots of the differences in the estimated Pearson correlation for $100,000$ randomly selected pairs of phecodes (out of all (${1,773\choose 2}$ pairs) comparing the ones obtained from synthetic data with those from MIMIC-IV for each of the 9 selected methods. The solid red line represents an exact match in correlation; the area within the red dotted lines indicates a correlation difference of less than 0.01.
For bivariate relationships preservation, Figure 34 shows the boxplots of the element-wise differences in the correlation matrix. Most methods overestimate correlations, except MedGAN, which underestimates them. This overestimation is even more pronounced in MIMIC-IV, where all methods, except MedGAN, consistently overestimate the correlations.
Fidelity Utility Privacy
MIMIC-III
MIMIC-IV
Table 2 presents the quantitative evaluation of key fidelity metrics. PBR outperforms other methods on most metrics, except for its higher CFD score (46.9), which reflects its inability to model bivariate correlations by design. CorGAN preserves distributional information best among the selected methods, particularly excelling in CFD due to its design for modeling associations.
When evaluated on the MIMIC-IV data, CorGAN outperforms the baselines in preserving the phecode-wise prevalence, achieving an MMD of 0.234 for absolute differences. However, PromptEHR performs best in relative difference metrics (RMSPE: 113.5; MAPE: 1418.7). Additionally, since synthetic data were trained on MIMIC-III, it is significantly easier to distinguish them from real MIMIC-IV data. Among all the selected methods, CorGAN remains the most effective, achieving lowest AUC of 0.927.
Utility
Obesity v.s. All Cancers Diabetes v.s. Hypertension
$\hat{\beta}$ 95% CI $\hat{\beta}$ 95% CI
Real Data MIMIC-III 0.871 (0.782 0.960) 1.219 (1.174 1.263)
MIMIC-IV 0.810 (0.782 0.838) 1.412 (1.388 1.437)
Synthetic Data Resample 0.868 (0.781 0.955) 1.194 (1.151 1.237)
Prevalence-based Random -0.125 (-0.316 0.066) 0.003 (-0.038 0.043)
Synthea 1.685 (1.614 1.756) 3.590 (3.512 3.668)
Plasmode -0.039 (-0.143 0.066) 0.027 (-0.118 0.171)
MedGAN 3.636 (3.453 3.819) 0.864 (0.822 0.905)
CorGAN 1.080 (0.970 1.191) 1.165 (1.120 1.209)
VAE - - 3.190 (3.063 3.318)
PromptEHR 0.514 (0.420 0.608) 0.597 (0.553 0.642)
EHRDiff 0.565 (0.527 0.603) 0.946 (0.891 1.000)
We first assess the analytical utility of the synthetic EHR data. Table 3 presents the logistic regression results. In the Obesity vs. All Cancers setting, the estimated log-odds ratios for the real data are 0.87 (95% CI: [0.78, 0.96]) in MIMIC-III and 0.81 (95% CI: [0.78, 0.84]) in MIMIC-IV. For Diabetes vs. Hypertension, the estimates are 1.22 (95% CI: [1.17, 1.26]) in MIMIC-III and 1.41 (95% CI: [1.39, 1.44]) in MIMIC-IV. Although the estimated directions of the two datasets are consistent, their magnitudes differ in the second task. For the synthetic data, PBR yields insignificant estimates for both tasks. Notably, CorGAN has estimated log-odds ratio of 1.08 (95% CI: [0.97, 1.19]) and 1.17 (95% CI: [1.12, 1.21]) in the two tasks, with confidence intervals substantially overlapping those of MIMIC-III in the second task. Notably, VAE encounters positivity issues in the first task, failing to generate obesity cases without any type of cancer, leading to failure in model convergence. This issue highlights its limitations in generating reliable synthetic data.
We then examine the predictive utility of synthetic EHR data (Table 2). In TSTR, all methods show performance degradation compared to TRTR on both datasets. MedGAN shows the smallest performance drop in AUC on both datasets, with a decrease of 0.063 on MIMIC-III and 0.076 on MIMIC-IV. In contrast, Plasmode and EHRDiff exhibit the most significant declines, with AUC decreases of 0.268 and 0.178, respectively, on MIMIC-III, and 0.218 for both on MIMIC-IV.
For TSRTR, on MIMIC-III, Plasmode becomes competitive with MedGAN, showing minor improvements in AUC (+0.0003) and ACC (+0.003). On MIMIC-IV, MedGAN, VAE, and PromptEHR perform competitively, each achieving a marginal ACC improvement of 0.004 in the downstream task.
Figure S7 and Table S2 presents results examining the performance in TSTR with varying prevalence of the phecodes. See Sections S1 and S2 for a more comprehensive discussion.
Privacy
For the MIR (Table 2), PBR and Synthea demonstrate the strongest performance on MIMIC-III, because these methods do not rely on training data to generate synthetic EHRs. Regarding the AIR, Synthea and Plasmode perform best, with predictive performance by attackers at 0.571 and 0.595, respectively. PBR follows closely, with a predictive performance of 0.622. Plasmode’s superior performance can be attributed to its dimension-by-dimension data generation approach, which omits the associations in the real dataset.
When evaluated on the MIMIC-IV dataset (assuming subjects in MIMIC-IV can be attacked), Plasmode and Synthea achieve the best results in lowering the risk of both membership and attribute inference attacks, mostly consistent with findings if MIMIC-III subjects are attacked.
A major contribution of this paper is to consolidate all seven synthetic data generation pipelines under one unified coding framework. To achieve this, we developed “SynthEHRella”, a Python package that integrates various generation approaches and evaluation metrics, facilitating future streamlined evaluation and methods selection. This also reduces the barrier to benchmarking multiple methods, making it a valuable tool for developing new approaches. The package is available at https://github.com/chenxran/synthEHRella.
4 DISCUSSION
This scoping review and benchmarking provide several key insights. First, PBR best resembles real MIMIC-III data although, by design, it does not capture bivariate associations. This finding indicates a major gap in synthetic data fidelity to be improved, and is consistent with previous work[yan2022multifaceted]. Second, CorGAN best preserves data distributions and analytical utility. However, it consistently underestimates prevalence (Figure S4). Third, Synthea offers strong privacy protection by not relying on training data but offers limited utility due to its restricted disease modules. Fourth, GAN-based methods are the most efficient, likely due to their lightweight, non-longitudinal design.
Based on our evaluation, we offer recommendations for practitioners to select synthetic EHR generation methods (Figure 5). When distributional shifts exist between training and testing samples, CorGAN is the preferred method unless downstream predictive performance is the priority, in which case MedGAN is the best choice. When the training and the testing populations are identical, we recommend MedGAN for downstream predictive tasks, CorGAN and PBR for better estimation of dimension-wise prevalence, and CorGAN for better estimation of associations among the phenotypes. Despite the inclusion of diverse model types, the top-performing methods are mostly GAN-based or PBR, which is consistent with previous benchmarks[yan2022multifaceted]. There are two possible explanations. First, methods like Synthea and PromptEHR generate data longitudinally, which may underperform on cross-sectional phenotype tasks. Second, emerging architectures (e.g., Transformer/Diffusion models) may not yet realize their full potential. However, these methods show promise for generating multimodal EHR data.
Figures S12S13 further examine trade-offs between fidelity/utility and privacy, and transportability across MIMIC-III and MIMIC-IV. The findings suggest that the trade-offs between non-privacy and privacy related metrics exist and are consistent across datasets, despite the systematic performance degradation caused by distributional differences. These align with results in the previous benchmarking study[yan2022multifaceted]. See Section S3 for detailed descriptions.
Our evaluation identifies several open opportunities. First, we showed that marginal PBR can generate synthetic data that resemble the real data, despite failing to capture correlations. However, even CorGAN with best fidelity performance, consistently underestimated these prevalence, leaving gaps for improvements. Second, as even the current most complicated methods did not consider the combination of both the structured (e.g., tabular code events and lab results) and unstructured (e.g., imaging, clinical notes) data, developing methods that account for the true multimodal complexity of EHR data, taking account of the temporal nature is a pressing open problem in synthetic EHR generation. Third, as from Figure 5, further investigation is required to exploit emerging architectures (e.g., Transformer/Diffusion models). Fourth, as generating synthetic multimodal EHR data with temporal features will become a common practice, new benchmarks and evaluation metrics are needed to assess such methods effectively. Finally, new methods are needed to incorporate formal privacy protections, e.g, differential privacy, to improve flexible control over the privacy-fidelity/utility balance in synthetic EHR generation. Other open opportunities are discussed in Section S4.
A major limitation of our study is our focus on only cross-sectional phenotype codes, excluding longitudinal evaluation of EHR data. We also did not evaluate conditional generation based on sociodemographic variables. Despite these gaps, our contribution is important for evaluating and inspiring further development and benchmarking of synthetic EHR methods. For example, our software platform SynthEHRella that implements all these generation methods can be extended to other data modalities and provide evaluation metrics in the expanded context. The remaining limitations of this paper are discussed in Section S4.
5 CONCLUSION
This paper presents a scoping review and a benchmarking of the existing representative methods with open-source codebases for generating synthetic EHR data using benchmark datasets. Under privacy concern, when distributional discrepancy exists between the training and testing samples, CorGAN is the best choice for dimension-wise prevalence estimation and downstream analytical tasks; MedGAN is recommended for training ML models for downstream use. Otherwise, when the distribution of the testing population is similar to that of the training data, MedGAN, PBR, and CorGAN are preferred choices for predictive modeling, dimension-wise prevalence estimation, and modeling associations among phenotypes, respectively. Despite the significant increase in the number of publications in the recent literature on synthetic EHR generation methods, we identify a few critical areas for future research. The benchmarking results highlight a substantial performance gap in fidelity between the existing methods and PBR, indicating the need for future methods research to narrow this gap. Finally, comprehensive benchmarking of the existing synthetic data generation methods that account for the longitudinal nature of EHR data and support conditional data generation is needed.
Funding Statement
This work was supported through grant DMS1712933 from the National Science Foundation and MI-CARES grant 1UG3CA267907 from the National Cancer Institute. The funders had no role in the design of the study; collection, analysis, or interpretation of the data; writing of the report; or the decision to submit the manuscript for publication.
Competing Interests Statement
The authors of this study have no competing interests to declare.
Contributorship Statement
The contributions of authors in this paper are summarized as follows: Xingran Chen (data curation, methodology, formal analysis, visualization, open-source package, writing), Zhenke Wu (formal analysis, methodology, supervision, writing), Xu Shi (critical revisions), Hyunghoon Cho (critical revisions), Bhramar Mukherjee (Conceptualization, methodology, supervision, writing, funding acquisition). All authors approved the final version of the manuscript.
Data Availability statement
The data used for benchmarking evaluation in this study were provided by the authors of MIMIC-III and MIMIC-IV and are available on PhysioNet upon request. The codebase used to implemented the benchmarking evaluation in this study is available at https://github.com/ chenxran/synthEHRella. There is no new dataset generated and released in this study.
Acknowledgment
Large language models (LLM)-assisted technologies, such as ChatGPT-4o, were used to improve the writing of this paper. The primary use of LLMs was to refine the language and grammar. All LLM suggestions were evaluated by the first author prior to their incorporation into the manuscript. LLMs were strictly limited to proposing edits to the initial drafts and were not used to generate any new materials.
[…]
Supplementary Materials
@@ Acknowledgment
Table S1: Summary statistics of the MIMIC-III and MIMIC-IV data. For age, gender, ethnicity, and diseases, the values represent percentages; counts are shown in parentheses.
@@
Figure S4: Boxplots of the ratio of the estimated PhecodeX-wise prevalence between the real and the synthetic data generated by CorGAN, grouped by PhecodeX categories defined in [10.1093/bioinformatics/btad655]. Y-axis represents the ratio of estimated prevalence, less than 1 means underestimation by the synthetic data. X-axis represents all the PhecodeX categories. Red line indicates the median of prevalence ratio of all phecodes. The significant over-estimation within the neonatal phenotype group in the MIMIC-IV dataset is because of its lack of neonatal population.
@@
(a) MIMIC-III.
(b) MIMIC-IV.
Figure S7: Scatterplots comparing the AUCTSTR and code frequency of selected phecodes in the MIMIC-III and MIMIC-IV datasets. Each scatterplot corresponds to a generation method and includes 20 phenotypes randomly selected from those with a prevalence greater than 0.1. Spearman’s $\rho$ correlation is reported for each plot. The top-left scatterplots display TRTR performance based on real data. Outlier phecodes in the TRTR evaluation from real data are annotated across all TSTR-versus-prevalence scatterplots for MIMIC-III and MIMIC-IV, respectively. Specifically, the outliers are Hypertension (CV_401), Disorders of fluid, electrolyte, and acid-base balance (EM_256), and Anemia (BI_164) in MIMIC-III; Anemia (BI_164), Fractures (MS_745), and Disorders of the thyroid gland (EM_200) in MIMIC-IV.
Table S2: AUCTSTR performance comparison of different methods for additional diseases in MIMIC-III and MIMIC-IV. PhecodeX codes, prevalence, and the AUCTRTR results of the selected diseases are reported in the table for reference.
MIMIC-III MIMIC-IV
Diseases Hypertension Diabetes Anxiety Depression Obesity Hypertension Diabetes Anxiety Depression Obesity
PhecodeX CV_401 EM_202 MB_288 MB_286 EM_236 CV_401 EM_202 MB_288 MB_286 EM_236
Prevalence (%) 47.6 25.4 11.0 16.0 5.4 46.0 21.7 13.0 20.8 12.6
TRTR Results 78.3 99.6 99.8 88.1 98.9 84.6 96.2 99.9 93.4 90.7
Resample 78.9 99.7 99.9 89.2 99.4 77.9 90.0 99.9 91.5 80.5
PBR 59.4 99.5 99.7 80.6 98.7 59.3 90.4 99.7 80.0 80.1
Synthea 63.1 67.9 53.1 53.1 54.9 72.6 75.8 64.8 57.1 65.0
Plasmode 48.8 97.6 91.4 81.6 49.3 48.5 90.9 93.9 81.4 50.1
MedGAN 71.9 98.8 98.4 84.6 95.5 77.0 92.4 98.9 89.7 75.7
CorGAN 70.3 99.3 99.6 83.9 97.7 71.4 89.0 99.6 85.7 77.2
VAE 69.5 97.4 66.3 58.7 52.1 72.3 88.3 82.9 69.3 34.0
PromptEHR 69.9 99.0 99.7 82.6 97.0 72.9 89.6 99.6 80.0 80.9
EHRDiff 61.1 99.1 98.2 79.4 96.8 62.9 89.6 99.5 79.7 77.8
Average 65.9 95.4 89.6 77.1 82.4 68.3 88.5 93.2 79.4 69.1
(a) Evaluated on MIMIC-III.
(b) Evaluated on MIMIC-IV.
Figure S8: Histogram of the Euclidean distances from each real EHR sample in MIMIC-III or MIMIC-IV datasets to its nearest synthetic EHR sample generated by included generation methods. The red solid line represents the sample mean, and the blue dotted line represents the sample median. Patients with no diagnosed phecode are excluded from the calculation.
(a) Evaluated on MIMIC-III.
(b) Evaluated on MIMIC-IV.
Figure S9: Cumulative distribution function (CDF) plots of the Euclidean distances from each real EHR sample in MIMIC-III or MIMIC-IV datasets to its nearest synthetic EHR sample generated by included generation methods. Higher curve values at a given distance cutoff indicate a greater proportion of real EHR samples with synthetic samples within that distance. Patients with no diagnosed phecodes are excluded from the calculation.
Table S3: Proportion of real EHR samples in MIMIC-III and MIMIC-IV that are perfectly matched with at least one synthetic EHR sample generated by selected methods. The values represent percentage (%). Patients with no diagnosed phecode are excluded from the calculation.
Method MIMIC-III MIMIC-IV
Resample 71.74 11.17
PBR 0.90 1.50
Synthea 0.78 5.52
Plasmode 1.56 0.02
MedGAN 13.28 6.65
Corgan 14.38 7.66
VAE 8.49 4.83
PromptEHR 1.28 0.06
EHRDiff 7.17 7.32
@@
Figure S10: The performance under varying sizes of the synthetic data ($M$) generated by CorGAN trained on full real data. The chosen evaluation metrics for the three perspectives are Maximum Mean Discrepancy (MMD), Train on Synthetic, Test on Real (TSTR) measured by Area Under the Curve (AUC), and F1 score for Attribute Inference Risk (AIR), respectively. The synthetic data generation process is repeated five times, and the mean evaluation metrics are reported. Note that for MMD and F1 score in AIR, lower values indicate better performance.
Figure S11: The performance of the synthetic data generated by CorGAN trained under varying sizes of the real data for training ($N$). The chosen evaluation metrics for the three perspectives are Maximum Mean Discrepancy (MMD), Train on Synthetic, Test on Real (TSTR) measured by Area Under the Curve (AUC), and F1 score for Attribute Inference Risk (AIR), respectively. The synthetic data generation process is repeated five times, and the mean evaluation metrics are reported. Note that for MMD and F1 score in AIR, lower values indicate better performance.
@@
Figure S12: Within-dataset Pearson correlation among different evaluation metrics in Table 2 for (a) MIMIC-III, and (b) MIMIC-IV, respectively. All “lower-is-better” evaluation metrics (e.g., Maximum Mean Discrepancy (MMD)) are flipped to their opposite numbers before calculating the Pearson correlation. Positive correlations (grids in red) indicate synergistic relationships between two metrics, and negative correlations (girds in blue) indicate trade-offs. No numbers are shown on the diagonal because here we focus on the relationships between distinct metrics evaluated on the same data; only lower triangular numbers are shown because of symmetry. For fidelity, Maximum Mean Discrepancy (MMD), Root Mean Squared Percentage Error (RMSPE), Correlation Frobenius Distance (CFD), and the Area Under the Curve (AUC) of discriminative prediction are included. For utility, the performance gap of AUC between Train on Synthetic, Test on Real (TSTR) and Train on Real, Test on Real (TRTR) are included, and similarly for the Train on Synthetic + Real, Test on Real (TSRTR) one. For privacy, the median of minimum Euclidean distance between each real medical record and the synthetic EHR dataset are included for Membership Inference Risk (MIR); the F1 score of predictive performance of 1-Nearest Neighbor (1-NN) are included for Attribute Inference Risk (AIR).